OpenStudy (anonymous):

derivative of log(sin^2(x))

OpenStudy (anonymous):

2cotx

OpenStudy (anonymous):

Thank you, but how do you get that?

OpenStudy (anonymous):

Chlorophyll's method.

OpenStudy (anonymous):

Got it, thanks guys :)

OpenStudy (raden):

using chain rule let u=sin^2(x) du/dx=2sin(x)cos(x) thus, y=log(u) dy/du=1/(u*ln(u)) = 1/(sin^2(x)*ln(sin^2(x)) y'=du/dx * dy/du = 2sin(x)cos(x) * 1/(sin^2(x)*ln(sin^2(x)) simplify it !

OpenStudy (raden):

i think ash.vasvani's and chlorophyll's answer still wrong because derivative of log(u) is not u'/u but should be 1/(u'*ln(u))

OpenStudy (raden):

am i right, @Chlorophyll ?

OpenStudy (anonymous):

I'm sorry to say you should double check the formula :)

OpenStudy (raden):

ok, nope... sometimes i do mistake too :D

OpenStudy (anonymous):

Yes, you do make mistake this time :) As most people know (ln x)' = x'/x = 1/x similarly, (ln u)' = u'/ u

OpenStudy (raden):

but, for this time u have mistaken again :P the original problem is y=log(...) not y=ln(...) hahahaha...

OpenStudy (anonymous):

I see what you mean now! Yes, I read the post so fast, however double check your the formula, still: y = log ( u ) --> y' = u' / u ln a

OpenStudy (raden):

yups, if like that i with u :) but a=base10, right ?

OpenStudy (anonymous):

Yep, to make it clear y = log ( u ) --> y' = u' / (u ln10)

OpenStudy (raden):

ok, thank u... nice discuss with u brother :)

OpenStudy (anonymous):

It would be nice to go to the bottom of discussion :)

OpenStudy (raden):

yea, but for this time i will go, i have a class know my student will angry to me if come late :)

OpenStudy (raden):

nothink, bye...